# Study of the decay pattern of $f_{0}(1370)$ as a $\kappa\bar{\kappa}$ molecular state

### v1

@@ III Numerical results and discussions
As we can see from Fig. 3(a) and (b), with the increase of $\sqrt{s}$ from $1.3$ GeV to $1.7$ GeV, the decay widths of the two-body channels such as $K\bar{K}$, $\pi\pi$ and $\eta\eta$ are stable. While for the four-body decay channels $4\pi$ and $K\bar{K}\pi\pi$, the decay width continuously increase along with the increase of $\sqrt{s}$. The dramatic increase in width of the $K\bar{K}\pi\pi$ channel is more evident in Fig. 3(b) than in Fig. 3(a) due to the scales. This phenomenon can be attributed to the increased phase space for the four-body decay and the threshold open for the $\kappa\bar{\kappa}$ and $\rho\rho$ intermediate states as $\sqrt{s}$ increases. The $\rho\rho$ threshold open effect is apparent as shown in Fig. 4, which will be discussed later. Due to such distortion by thresholds, the resonance will manifest with different lineshapes in different decay spectra, just as discussed in [66]. In other point of view, the analytic continuation of a system with pole deep inside the complex plane can be highly non-trivial, and the projection onto the real axis, where experiments are performed, can differ widely depending on the complexity of the system. In Fig. 3 we can see that before energy point about $1.45$ GeV, the largest contribution to the total width comes from the $K\bar{K}$ channel, the order of these partial widths is $\Gamma_{K\bar{K}}>\Gamma_{4\pi}>\Gamma_{\pi\pi}>\Gamma_{K\bar{K}\pi\pi}>\Gamma_{\eta\eta}$. When $\sqrt{s}$ is larger than $1.45$ GeV, the partial width of the $4\pi$ channel exceed that of the $K\bar{K}$ channel, becoming the most dominant contribution. However, when $\sqrt{s}$ is less than about $1.34$ GeV, the partial width of $4\pi$ channel is even smaller than that of the $\pi\pi$ channel. Similarly, along with the increase of $\sqrt{s}$, the large difference between the partial widths of the $K\bar{K}\pi\pi$ and $\pi\pi$ channels at small $\sqrt{s}$ start to reduce gradually.
@@
Analysis on the $2\pi^{+}2\pi^{-}$ spectrum in [21] gives the ratio of $Br(\rho^{0}\rho^{0}\to 2\pi^{+}2\pi^{-}):Br(\sigma\sigma\to 2\pi^{+}2\pi^{-})$ as $3:7$. When comparing this value to our result shown in Table 3, which is approximately $1:4$, it implies that there might be an overestimation of the contributions from the $\sigma\sigma$ intermediate state and an underestimation of the contributions from the $\rho\rho$ intermediate state in our result. However, remind that our result is obtained at fixed energy $\sqrt{s}=1.37$ GeV, and the branching ratio are obtained by integrating over the mass distribution of $f_{0}(1370)$. Furthermore, this ratio is rather sensitive to the parametrization of the $\sigma$, which we adopted is different from that in [21].
@@
Figure 3: The dependence of the partial widths of different channels on the initial energy $\sqrt{s}$ with fixed initial coupling $g_{F\kappa\bar{\kappa}}=35$ GeV and cut-off parameter $\alpha=2$. Because the scales in widths for some channels are much smaller than others, we present them in two separate figures with distinct scale for clarity. (a) $K\bar{K}$, $\pi\pi$, $4\pi$ and $K\bar{K}\pi\pi$ channels. Different specific $4\pi$ states are also presented. (b) $K\bar{K}\pi\pi$ and $\eta\eta$ channels. Different specific $K\bar{K}\pi\pi$ states are also presented. Note that $\Gamma(K^{+}K^{-}\pi^{+}\pi^{-})=\Gamma(K^{0}\bar{K}^{0}\pi^{+}\pi^{-})$, $\Gamma(K^{+}\bar{K}^{0}\pi^{0}\pi^{-})=\Gamma(K^{0}K^{-}\pi^{+}\pi^{0})$ and $\Gamma(K^{+}K^{-}\pi^{0}\pi^{0})=\Gamma(K^{0}\bar{K}^{0}\pi^{0}\pi^{0})$. For the $\eta\eta$ channel, $\Gamma(F\to\eta\eta)$ is presented.
@@
For a more reliable comparison, we plot the respective partial widths of these two parts varying along with the initial energy $\sqrt{s}$ in Fig. 4, to see when integrating over the mass distribution what tendency this ratio will show. Fig. 4(a) and (b) refer to the $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ states, respectively. The black solid line indicates the total width of the specific $4\pi$ state, which includes the interference between the $\rho\rho$ and $\sigma\sigma$ intermediate states. The red long dashed line and the blue short dashed line indicate the exclusive widths from the $\sigma\sigma$ and $\rho\rho$ intermediate states respectively. In Fig. 4(a), we observe that the increase in the $\sigma\sigma\to 4\pi$ process is moderate, whereas the $\rho\rho\to 4\pi$ process experiences a rapid increase once the energy $\sqrt{s}$ surpass the $\rho\rho$ threshold, leading to a gradually reduced discrepancy between the $\sigma\sigma$ and $\rho\rho$ cases. Therefore, when considering the integration over the whole mass distribution, the proportion of the contributions from the $\rho\rho$ intermediate state might be improved.
@@
Figure 4: The dependence of the partial widths of $4\pi$ states on the initial energy $\sqrt{s}$. Contributions from different intermediate channels $\sigma\sigma$ and $\rho\rho$ are also plotted. (a) For the $2\pi^{+}2\pi^{-}$ states. (b) For the $\pi^{+}\pi^{-}2\pi^{0}$ states.
@@
Such a seeming underestimation of the theoretical $\rho\rho$ contributions might be some problems in experimental analyses. The model used in experiments might overestimate the contributions of the $f_{0}(1370)$ decay to $4\pi$ through the $\rho\rho$ channel, as some of the $\rho\rho\to 4\pi$ events that originate from the $f_{0}(1500)$ decay may be mistakenly attributed to the $f_{0}(1370)$ decay. It is possible as $f_{0}(1500)$ is very close in mass to $f_{0}(1370)$ and is suggested to be a $\rho\rho$ molecular state [18]. If so, this will consequently lead to an overestimation of the branching ratio of $f_{0}(1370)$ decay to $4\pi$.
@@
We also plot the direct sum of the exclusive widths of the $\rho\rho$ and $\sigma\sigma$ cases, which is indicated by green dashed line and marked as “DirectSum” in Fig. 4. We found that it is almost coincides with the black solid line along all the $\sqrt{s}$ for both the $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ channels, which implies that the $\rho\rho$ and $\sigma\sigma$ channels are almost orthogonal and their interference is negligible. This is because the $\rho\pi\pi$ vertex is of $P$-wave while the $\sigma\pi\pi$ vertex is of $S$-wave. Such a phenomenon is also confirmed in the analysis in Ref. [70].
@@
Table 4: Collection of the available experiments and our results in the case of $\alpha=1/\mathbf{2}/3$. The initial coupling $g_{F\kappa\bar{\kappa}}$ is fixed to be $35$ GeV.
@@
In Table 4, the relation between the $K\bar{K}$ and $\pi\pi$ channels has two kinds of results. Qualitatively, measurements from the $\pi\pi$ and $\pi p$ scattering obtained at the mass of $f_{0}(1370)$ [71] support that $\Gamma(F\to K\bar{K})$ is larger than $\Gamma(F\to\pi\pi)$, which is consistent with our results at $\sqrt{s}=1.37$ GeV. In contrast, measurements from the $J/\psi$ hadronic decay [52], $J/\psi$ radiative decay [48], the $pp$ central production [41] and the $p\bar{p}$ annihilation at rest [72] show an inverse branching fraction ratio. As mentioned in the Introduction, the data of $J/\psi$ hadronic decay recoilling $\phi$ even gives a very small value for this ratio of $0.08\pm 0.08$ due to the conspicuous signal in $\pi\pi$ but absence of any corresponding peak in $K^{+}K^{-}$. We attempt to explain this by the destructive interference between the two parts in the $f_{0}(1370)$ wavefunction: $s\bar{s}$ and the molecular component $\kappa\bar{\kappa}$. Although the $s\bar{s}$ component is not included in the wavefunction of Eq. (3) due to the approximated $X=1$ is taken in our calculation above. However, there would not be a pure molecular state, the wavefunction could include a small component of $\bar{s}. ss$$\bar{s}$ is produced in $J/\psi$ decay recoiling $\phi$ as leading order and it strongly couple to $K\bar{K}$ but contribute to $\pi\pi$ with a more weak strength, leading to the destructive effect in the $K\bar{K}$ channel but not affecting the $\pi\pi$ channel. Such an $s\bar{s}$ desctructive effect might also appear in the $pp$ central production and $J/\psi$ radiative decay but appear in the $\pi p$ scattering with a smaller fraction. The more quantitative discussion about this issue is complex and needs more reliable analyses on experiments.
@@
The analysis in Ref. [76] carefully considered the Bose symmetry interference effects and found that these effects cause the ratios between different specific $4\pi$ states to obviously differ from the naive counting values without considering these effects, as what is evaluated in [21] based on the $\text{Br}(2\pi^{+}2\pi^{-})$ to obtain these ratios. Hence, the estimated ratios of $Br(2\pi^{+}2\pi^{-}):Br(\pi^{+}\pi^{-}2\pi^{0}):Br(4\pi^{0})$ in [21] is not reliable. While such effects have been fully considered in our calculations.
@@ IV Summary
In summary, motivated by the phenomenology that the $f_{0}(980)$, $f_{0}(1370)$ and $f_{0}(1790)$ appear successively in the $\pi\pi$ spectrum in $J/\psi$ hadronic decay recoiling $\phi$, where the $s\bar{s}$ is produced in leading order, as well as considering the $f_{0}(980)$ and $f_{0}(1790)$ are plausible $K\bar{K}$ and $K^{*}\bar{K}^{*}$ molecules respectively, we assume that $f_{0}(1370)$ is a $\kappa\bar{\kappa}$ molecular state and calculate the partial widths of various channels in $f_{0}(1370)$ decay, including $K\bar{K}$, $\pi\pi$, $\eta\eta$, $4\pi$ and $K\bar{K}\pi\pi$.
@@
We first use the initial coupling $g_{F\kappa\bar{\kappa}}$ ($F$ indicates $f_{0}(1370)$) determined by the Weinberg criterion with complex $\kappa$ mass, and found that the total width of $f_{0}(1370)$ is much smaller than its measured widths with a reasonable form factor. It is reasonable that the Weinberg criterion derived from bound state with narrow components does not suit our case with broad components $\kappa$. Then, we fit the width of $f_{0}(1370)$ listed in the RPP by taking the initial coupling $g_{F\kappa\bar{\kappa}}$ as fitting parameter and found that this coupling should be in the range of $25\sim 40$ GeV with cut-off parameter $\alpha=2\sim 3$. But the decay pattern is independent of this initial coupling because it is a common coupling for all the decay channels.
@@
At the nominal mass of $f_{0}(1370)$: $m_{F}=1.37$ GeV and with $g_{F\kappa\bar{\kappa}}=35$ GeV and $\alpha=2$, the $K\bar{K}$ channel has the largest partial width, followed by the $\pi\pi$ and $4\pi$ channels. The $\pi\pi$ and $4\pi$ channels have comparable partial widths. These three channels are the most dominant decay channels in $f_{0}(1370)$ decay. Different strategies for determining $\kappa K\eta$ coupling result in significant variations in the partial width of the $\eta\eta$ channel. We have found the distinct decay properties in $\eta\eta$ channel between $\kappa\bar{\kappa}$ and $K^{*}\bar{K^{*}}$ molecules, which can be examined in the BESIII data of $J/\psi\to\phi\eta\eta$. The partial width of the $K\bar{K}\pi\pi$ channel is about $300$ keV in total. Notice that although the $K\bar{K}\pi\pi$ channel proceeds through tree-level decay, while it has very limited four body phase space around $1.37$ GeV. Hence, the major contributions to the total width of $f_{0}(1370)$ are from the $K\bar{K}$, $\pi\pi$ and $4\pi$ channels.
@@
We also test the dependence of the partial widths of different channels on the initial energy $\sqrt{s}$. The decay widths of the two-body channels such as $K\bar{K}$, $\pi\pi$ and $\eta\eta$ are stable with the variation of $\sqrt{s}$. While the decay widths of the four-body decay channels $4\pi$ and $K\bar{K}\pi\pi$ continuously increase along with the increase of $\sqrt{s}$. With $\sqrt{s}$ larger than $1.45$ GeV, the partial width of the $4\pi$ channel exceed that of the $K\bar{K}$ channel, becoming the most dominant decay channel. Such phenomenon might cause different manifestation of $f_{0}(1370)$ in different spectra.
@@
Not all of the ratios reflecting the decay pattern in our calculations are consistent with the current data, such as the $4\pi$ dominant conclusion and the ratio of $K\bar{K}$ over $\pi\pi$ in some measurements. However, in $f_{0}(1370)$ decay, most of the current data are model-dependent, relying on fitting sums of Breit-Wigner functions and no single experimental number on branching ratios nor ratios thereof are determined because they are conflicting. Hence, these inconsistencies could not rule out the $\kappa\bar{\kappa}$ assignment for $f_{0}(1370)$. The low $K\bar{K}$ branching ratio contradict with our results mostly due to the absence of corresponding peak in $K\bar{K}$ spectrum, such as in $J/\psi$ hadronic decay recoiling $\phi$. We attempt to explain that by the destructive interference in decay to $K\bar{K}$ between the $s\bar{s}$ and $\kappa\bar{\kappa}$ in the $f_{0}(1370)$ wavefunction. The $4\pi$ spectra is also elusive because it is hard to disentangle the contribution from $f_{0}(1370)$ and $f_{0}(1500)$ due to the strongly $s$-dependent width of $f_{0}(1370)$. The ratios between different $4\pi$ states can be related to decay mechanisms as the intermediate states $\rho\rho$ and $\sigma\sigma$ have distinct features in $4\pi$ decay. The discrepancy in the $\rho\rho$ contributions between our results and data might be attributed to the overestimation in experiments for mistakenly counting the $\rho\rho\to 4\pi$ events from the $f_{0}(1500)$ decay.
@@
We need further reliable analyses on the $f_{0}(1370)$ in both theoretical and experimental aspects. The BESIII Collaboration has recorded data with significantly improved quality and statistics. They can attempt to research the $f_{0}(1370)$ signal in the $K\bar{K}\pi\pi$ channel. Such measurements can help us to test the $\kappa\bar{\kappa}$ molecule assignment for $f_{0}(1370)$ and the pattern for the different specific $K\bar{K}\pi\pi$ states also has unique feature in this scenario, which can be used as a further check.

### v2

@@ III Numerical results and discussions
As we can see from Fig. 3(a) and (b), with the increase of $\sqrt{s}$ from $1.3$ GeV to $1.7$ GeV, the decay widths of the two-body channels such as $K\bar{K}$, $\pi\pi$ and $\eta\eta$ are stable. While for the four-body decay channels $4\pi$ and $K\bar{K}\pi\pi$, the decay width continuously increase along with the increase of $\sqrt{s}$. The dramatic increase in width of the $K\bar{K}\pi\pi$ channel is more evident in Fig. 3(b) than in Fig. 3(a) due to the scales. This phenomenon can be attributed to the increased phase space for the four-body decay and the threshold open for the $\kappa\bar{\kappa}$ and $\rho\rho$ intermediate states as $\sqrt{s}$ increases. The $\rho\rho$ threshold open effect is apparent as shown in Fig. 4, which will be discussed later. As a consequence of these threshold effects, the apparent lineshape of the resonance varies depending on the specific decay channel, consistent with the analysis presented in [66]. From another perspective, the analytic continuation of a system with pole deep inside the complex plane can be highly non-trivial, and the projection onto the real axis, where experiments are performed, can differ widely depending on the complexity of the system. In Fig. 3 we can see that before energy point about $1.45$ GeV, the largest contribution to the total width is from the $K\bar{K}$ channel, the order of these partial widths is $\Gamma_{K\bar{K}}>\Gamma_{4\pi}>\Gamma_{\pi\pi}>\Gamma_{\eta\eta}>\Gamma_{K\bar{K}\pi\pi}$. When $\sqrt{s}$ is larger than $1.45$ GeV, the partial width of the $4\pi$ channel exceeds that of the $K\bar{K}$ channel, becoming the most dominant contribution. However, when $\sqrt{s}$ is less than about $1.34$ GeV, the partial width of $4\pi$ channel is even smaller than that of the $\pi\pi$ channel. Similarly, along with the increase of $\sqrt{s}$, the large difference between the partial widths of the $K\bar{K}\pi\pi$ and $\pi\pi$ channels at small $\sqrt{s}$ start to reduce gradually.
@@
Analysis on the $2\pi^{+}2\pi^{-}$ spectrum in [21] gives the ratio of $Br(\rho^{0}\rho^{0}\to 2\pi^{+}2\pi^{-}):Br(\sigma\sigma\to 2\pi^{+}2\pi^{-})$ as $3:7$. When comparing this value to our result shown in Table 3, which is approximately $1:4$, it implies that there might be an overestimation of the contributions from the $\sigma\sigma$ intermediate state and an underestimation of the contributions from the $\rho\rho$ intermediate state in our result. However, remind that our result is obtained at fixed energy $\sqrt{s}=1.37$ GeV, and the branching ratios are obtained by integrating over the mass distribution of $f_{0}(1370)$. Furthermore, this ratio is rather sensitive to the parametrization of the $\sigma$, which we adopted is different from that in [21].
@@
Figure 3: The dependence of the partial widths of different channels on the center-of-mass energy $\sqrt{s}$ with fixed coupling $g_{F\kappa\bar{\kappa}}=35$ GeV and cut-off parameter $\alpha=2$. Because the scales in widths for some channels are much smaller than others, we present them in two separate figures with distinct scale for clarity. (a) $K\bar{K}$, $\pi\pi$, $4\pi$ and $K\bar{K}\pi\pi$ channels. Different specific $4\pi$ states are also presented. (b) $K\bar{K}\pi\pi$ and $\eta\eta$ channels. Different specific $K\bar{K}\pi\pi$ states are also presented. Note that $\Gamma(K^{+}K^{-}\pi^{+}\pi^{-})=\Gamma(K^{0}\bar{K}^{0}\pi^{+}\pi^{-})$, $\Gamma(K^{+}\bar{K}^{0}\pi^{0}\pi^{-})=\Gamma(K^{0}K^{-}\pi^{+}\pi^{0})$ and $\Gamma(K^{+}K^{-}\pi^{0}\pi^{0})=\Gamma(K^{0}\bar{K}^{0}\pi^{0}\pi^{0})$. For the $\eta\eta$ channel, $\Gamma(F\to\eta\eta)$ is presented.
@@
For a more reliable comparison, we plot the respective partial widths of these two parts varying along with the centter-of-mass energy $\sqrt{s}$ in Fig. 4, to investigate the resulting trend of their ratio after performing the mass distribution integration. Fig. 4(a) and (b) refer to the $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ states, respectively. The black solid line indicates the total width of the specific $4\pi$ state, which includes the interference between the $\rho\rho$ and $\sigma\sigma$ intermediate states. The red long dashed line and the blue short dashed line indicate the exclusive widths from the $\sigma\sigma$ and $\rho\rho$ intermediate states respectively. In Fig. 4(a), we observe that the increase in the $\sigma\sigma\to 4\pi$ process is moderate, whereas the $\rho\rho\to 4\pi$ process experiences a rapid increase once the energy $\sqrt{s}$ surpass the $\rho\rho$ threshold, leading to a gradually reduced discrepancy between the $\sigma\sigma$ and $\rho\rho$ cases. Therefore, when considering the integration over the whole mass distribution, the proportion of the contributions from the $\rho\rho$ intermediate state might be improved.
@@
Figure 4: The dependence of the partial widths of $4\pi$ states on the center-of-mass energy $\sqrt{s}$. Contributions from different intermediate channels $\sigma\sigma$ and $\rho\rho$ are also plotted. (a) For the $2\pi^{+}2\pi^{-}$ states. (b) For the $\pi^{+}\pi^{-}2\pi^{0}$ states.
@@
Such an apparent underestimation of the theoretical $\rho\rho$ contributions might point to issues in experimental analyses. The model used in experiments might overestimate the contributions of the $f_{0}(1370)$ decay to $4\pi$ through the $\rho\rho$ channel, as some of the $\rho\rho\to 4\pi$ events that originate from the $f_{0}(1500)$ decay may be mistakenly attributed to the $f_{0}(1370)$ decay. It is possible as $f_{0}(1500)$ is very close in mass to $f_{0}(1370)$ and is suggested to be a $\rho\rho$ molecular state [18]. If so, this will consequently lead to an overestimation of the branching ratio of $f_{0}(1370)$ decay to $4\pi$.
@@
We also plot the direct sum of the exclusive widths of the $\rho\rho$ and $\sigma\sigma$ cases, which is indicated by green dashed line and marked as “DirectSum” in Fig. 4. We find that it is almost coincides with the black solid line along all the $\sqrt{s}$ for both the $2\pi^{+}2\pi^{-}$ and $\pi^{+}\pi^{-}2\pi^{0}$ channels, which implies that the $\rho\rho$ and $\sigma\sigma$ channels are almost orthogonal and their interference is negligible. This is because the $\rho\pi\pi$ vertex is of $P$-wave while the $\sigma\pi\pi$ vertex is of $S$-wave. Such a phenomenon is also confirmed in the analysis in Ref. [70].
@@
Table 4: Collection of the available experiments and our results in the case of $\alpha=1/\mathbf{2}/3$. The coupling $g_{F\kappa\bar{\kappa}}$ is fixed to be $35$ GeV.
@@
In Table 4, the relation between the $K\bar{K}$ and $\pi\pi$ channels has two kinds of results. Qualitatively, measurements from the $\pi\pi$ and $\pi p$ scattering obtained at the mass of $f_{0}(1370)$ [71] support that $\Gamma(F\to K\bar{K})$ is larger than $\Gamma(F\to\pi\pi)$, which is consistent with our results at $\sqrt{s}=1.37$ GeV. In contrast, measurements from the $J/\psi$ hadronic decay [52], $J/\psi$ radiative decay [48], the $pp$ central production [41] and the $p\bar{p}$ annihilation at rest [72] show an inverse branching fraction ratio. As mentioned in the Introduction, the data of $J/\psi$ hadronic decay recoiling against $\phi$ even gives a very small value for this ratio of $0.08\pm 0.08$ due to the conspicuous signal in $\pi\pi$ but absence of any corresponding peak in $K^{+}K^{-}$. We attempt to explain this by the destructive interference between the two parts in the $f_{0}(1370)$ wavefunction: $s\bar{s}$ and the molecular component $\kappa\bar{\kappa}$. Although the $s\bar{s}$ component is not included in the wavefunction of Eq. (3) because the approximated $X=1$ is taken in our calculation above. However, there would not be a pure molecular state, the $f_{0}(1370)$ wavefunction could include a small component of $\bar{s}. ss$$\bar{s}$ is produced in $J/\psi$ decay recoiling against $\phi$ as leading order and it strongly couple to $K\bar{K}$ but contribute to $\pi\pi$ with significantly weaker strength, leading to a possible destructive effect in the $K\bar{K}$ channel but not affecting the $\pi\pi$ channel. Such an $s\bar{s}$ destructive effect might also appear in the $pp$ central production and $J/\psi$ radiative decay but appear in the $\pi p$ scattering with a smaller fraction. The more quantitative discussion about this issue is complex and needs more reliable analyses on experiments.
@@
The analysis in Ref. [76] carefully considered the Bose symmetry interference effects and found that these effects cause the ratios between different specific $4\pi$ states to obviously differ from the naive counting values without considering these effects, as what is evaluated in [21] based on the $Br(2\pi^{+}2\pi^{-})$ to obtain these ratios. Hence, the estimated ratios of $Br(2\pi^{+}2\pi^{-}):Br(\pi^{+}\pi^{-}2\pi^{0}):Br(4\pi^{0})$ in [21] is not reliable. While such effects have been fully considered in our calculations.
@@ IV Summary
In summary, motivated by the fact that the $f_{0}(980)$ and $f_{0}(1790)$ are respectively plausible $K\bar{K}$ and $K^{*}\bar{K}^{*}$ molecules and the phenomenology that the $f_{0}(980)$, $f_{0}(1370)$ and $f_{0}(1790)$ appear successively in the $\pi\pi$ spectrum in $J/\psi$ hadronic decay recoiling against $\phi$, where the $s\bar{s}$ is produced in leading order, we assume that $f_{0}(1370)$ is a $\kappa\bar{\kappa}$ molecular state and calculate the partial widths of various channels in $f_{0}(1370)$ decay, including $K\bar{K}$, $\pi\pi$, $\eta\eta$, $4\pi$ and $K\bar{K}\pi\pi$.
@@
We first use the coupling $g_{F\kappa\bar{\kappa}}$ ($F$ indicates $f_{0}(1370)$) determined by the Weinberg criterion with complex $\kappa$ mass, and find that the total width of $f_{0}(1370)$ is much smaller than its measured widths with a reasonable form factor. It is reasonable that the Weinberg criterion derived from bound state with narrow components does not suit our case with broad components $\kappa$. Then, we fit the width of $f_{0}(1370)$ listed in the RPP by taking the coupling $g_{F\kappa\bar{\kappa}}$ as fitting parameter and find that this coupling should be in the range of $25\sim 40$ GeV with cut-off parameter $\alpha=2\sim 3$. But the decay pattern is independent of this coupling because it is a common coupling for all the decay channels.
@@
At the nominal mass of $f_{0}(1370)$: $m_{F}=1.37$ GeV and with $g_{F\kappa\bar{\kappa}}=35$ GeV and $\alpha=2$, the $K\bar{K}$ channel has the largest partial width, followed by the $4\pi$ and $\pi\pi$ channels. The $\pi\pi$ and $4\pi$ channels have comparable partial widths. These three channels are the most dominant decay channels in $f_{0}(1370)$ decay. Different strategies for determining $\kappa K\eta$ coupling result in significant variations in the partial width of the $\eta\eta$ channel. We have found the distinct decay properties in $\eta\eta$ channel between $\kappa\bar{\kappa}$ and $K^{*}\bar{K^{*}}$ molecules, which can be examined in the BESIII data of $J/\psi\to\phi\eta\eta$. The partial width of the $K\bar{K}\pi\pi$ channel is about $300$ keV in total. Notice that although the $K\bar{K}\pi\pi$ channel proceeds through tree-level decay, while it has very limited four-body phase space around $1.37$ GeV.
@@
We also test the dependence of the partial widths of different channels on the center-of-mass energy $\sqrt{s}$. The decay widths of the two-body channels such as $K\bar{K}$, $\pi\pi$ and $\eta\eta$ are stable with the variation of $\sqrt{s}$. While the decay widths of the four-body decay channels $4\pi$ and $K\bar{K}\pi\pi$ continuously increase along with the increase of $\sqrt{s}$. With $\sqrt{s}$ larger than $1.45$ GeV, the partial width of the $4\pi$ channel exceed that of the $K\bar{K}$ channel, becoming the most dominant decay channel. Such phenomenon might cause different manifestation of $f_{0}(1370)$ in different spectra.
@@
Not all of the ratios reflecting the decay pattern in our calculations are consistent with the current data, such as the $4\pi$ dominant conclusion and the ratio of $K\bar{K}$ over $\pi\pi$ in some measurements. However, in $f_{0}(1370)$ decay, most of the current data are model-dependent, relying on fitting sums of Breit-Wigner functions and no single experimental number on branching ratios nor ratios thereof are determined because they are conflicting. Hence, these inconsistencies can not rule out the $\kappa\bar{\kappa}$ assignment for $f_{0}(1370)$. The small $K\bar{K}$ branching ratio due to the absence of corresponding peak in $K\bar{K}$ spectrum, such as in $J/\psi$ hadronic decay recoiling against $\phi$, contradict with our result. We attempt to explain that by the destructive interference in decay to $K\bar{K}$ between the $s\bar{s}$ and $\kappa\bar{\kappa}$ components in the $f_{0}(1370)$ wavefunction. The $4\pi$ spectra is also elusive because it is hard to disentangle the contribution from $f_{0}(1370)$ and $f_{0}(1500)$ due to the strongly $s$-dependent width of $f_{0}(1370)$. The ratios between different $4\pi$ states can be related to decay mechanisms as the intermediate states $\rho\rho$ and $\sigma\sigma$ have distinct features in $4\pi$ decay. The discrepancy in the $\rho\rho$ contributions between our results and data might be attributed to the overestimation in experiments for mistakenly counting the $\rho\rho\to 4\pi$ events from the $f_{0}(1500)$ decay.
@@
We need further reliable analyses on the $f_{0}(1370)$ in both theoretical and experimental aspects. The BESIII Collaboration has recorded data with significantly improved quality and statistics. They can attempt to search for the $f_{0}(1370)$ signal in the $K\bar{K}\pi\pi$ channel. Such measurements can help us to test the $\kappa\bar{\kappa}$ molecule assignment for $f_{0}(1370)$. The pattern for the different specific $K\bar{K}\pi\pi$ states also has unique feature in this scenario, which can be used as a further check.